OpenStudy (anonymous):

find the integration of sin^2 x dx=???

OpenStudy (anonymous):

Let, I = int [(x^2)sinx dx] u = x^2, du = 2x dx dv = sinx dx, v = -cosx I = -(x^2)cosx - int [-2xcosx dx] I = -(x^2)cosx + int [2xcosx dx] Again integrate the second part by parts, int [2xcosx dx] u = 2x du = 2 dx dv = cosx dx v = sinx int [2xcosx dx] = 2xsinx - int [2sinx dx] int [2xcosx dx] = 2xsinx + 2cosx +c I = -(x^2)cosx + 2xsinx + 2cosx + c

OpenStudy (anonymous):

∫ sin²x dx since cos2x = 1 - 2sin²x sin²x = ½ - ½cos2x ∫( ½ - ½cos2x ) = ½x -¼sin2x + C @qamy

OpenStudy (anonymous):

@qamy Your integration is complete.

OpenStudy (anonymous):